If the line passing through the points (a,1) and (4,7) is parrallel to the line passing through the points (3,10) and (a+4,3) what is the value of a?
Answer by stanbon(75887)

If the line passing through the points (a,1) and (4,7) is parallel to the line passing through the points (3,10) and (a+4,3) what is the value of a?
----------------------
(7-1)/(4-a) = (10-30)/(3-(a+4))
----
6/(4-a) = -20/(-a-1)
----
6/(4-a) = 20/(a+1)
Cross multiply and solve for "a":
6a+5 = 80-20a
26a = 75
a = 2.8846...
Cheers,
Stan H.